# A new class of general linear method with inherent quadratic stability for solving stiff differential systems

### v1

@@ 4.1. Van der Pol problem
Here, the initial conditions are taken as $y(0)=2$ and $z(0)=-\frac{2}{3}+\frac{10}{81}\epsilon-\frac{292}{2187}\epsilon^{2}-\frac{1814}{19683}\epsilon^{3}+\mathcal{O}(\epsilon^{4})$. We have solved this problem for $\epsilon=10^{-6}$ and the interval $t\in[0,0.5]$, with end point $T=0.5$.
@@ 4.2. Burgers’ Equation
Where $d>0$ is the coefficient of viscosity (diffusion coefficient), which can often become very large. Burgers’ equation can be used to model a lot of different problems [1, 9]. For instance, Burgers’ equation is another way to look at the Navier-Stokes equations. It has nonlinear parts like unknown functions multiplied by a first derivative and higher-order parts multiplied by a small constant.
For illustration purposes, we choose the final space point $L=1$ and the final time point $T=1$. The spatial discretization is done by taking step size k=1/(M1)k=1/(M-1), where $M$ is the number of spatial grid points, for $x\in[0,1]$ where $x_{0}=0\text{ and }L=1$. We use the central difference approximation to discretize the second-order spatial derivative (diffusion term) on the right-hand side of (4.5) and the upwind scheme to discretize the first-order derivative (convection term) [1].
@@ 4.3. Gray-Scott Model
$\nabla^{2}\phi_{i,j}\approx\frac{\phi_{i+1,j}+\phi_{i-1,j}+\phi_{i,j+1}+\phi_{i,j-1}-4\phi_{i,j}}{k^{2}},$ (4.9)
@@
where $\Omega_{c}$ denotes a small square centered in $\Omega$.
@@ 5. Results and Discussion
The Gray–Scott reaction–diffusion system provides a challenging test of robustness in PDEs with complex dynamics. Convergence and order are validated and are reported in Table 3 and Figure 6. The proposed integrators of order $p=1,3$ produce less error than the $p=s$ class of [14], and the method of order $p=2,4$ is competitive for the choice of parameters made while constructing the example method. A different choice of free parameter, along with a suitable minimization, would lead to a more efficient class with the same or better efficiency.
@@
  • Also, the order estimation given in the Figures 1, 3, and 4 confirms that our proposed classes are free from any kind of order reduction.
  • The numerical solution at $t=0.5$ for different diffusion coefficients, as illustrated in Figure 4, preserves the parabolic character of the problem.
(a) $k=1/9$
(b) $k=1/19$
Figure 6. Convergence plot for Gray-Scott model
@@ 6. Conclusion
We have explained the construction of implicit GLMQS and presented four classes of $A$-stable and $L$-stable implicit numerical solvers for solving time-dependent stiff differential equations. The proposed classes of GLM are based on Nordsieck input vectors that satisfy the inherent quadratic stability criteria, along with the minimization of the error constant. Example methods of order $p=1,2,3,$ and $4$ are derived, and their governing coefficient matrices are reported in Section 3. We have also provided three numerical illustrations of time-dependent IVPs, both ODEs and PDEs, to test our proposed classes. The numerical results reported for all the van der Pol problem (ODE) and for both PDEs, the Burgers’ equation and Gray-Scott model, show that our methods are convergent and achieve the expected order of accuracy. In future articles, we will be focusing on the construction of higher-order methods and the development of implicit-explicit numerical schemes using the present classes of implicit schemes for solving time-dependent partitioned differential equations.

### v2

@@ 4.1. Van der Pol problem
Here, the initial conditions are taken as $y(0)=2$ and $z(0)=-\frac{2}{3}+\frac{10}{81}\epsilon-\frac{292}{2187}\epsilon^{2}-\frac{1814}{19683}\epsilon^{3}+\mathcal{O}(\epsilon^{4})$. We have solved this problem for $\epsilon=10^{-6}$, for which the problem becomes highly stiff in nature. The interval of integration $t\in[0,0.5]$, with end point $T=0.5$.
@@ 4.2. Burgers’ Equation
Here $d>0$ is the coefficient of viscosity (diffusion coefficient), which can often become very large. Burgers’ equation can be used to model a lot of different problems [1, 9]. For instance, Burgers’ equation is another way to look at the Navier-Stokes equations. It has nonlinear parts like unknown functions multiplied by a first derivative and higher-order parts multiplied by a small constant.
For illustration purposes, we choose the final space point $L=1$ and the final time point $T=1$. The spatial discretization is done by taking step size k=1/(M1)k=1/(M-1), where $M$ is the number of spatial grid points, for $x\in[0,1]$ where $x_{0}=0\text{ and }L=1$. We use the central difference approximation to discretize the second-order spatial derivative (diffusion term) on the right-hand side of (4.5) and the upwind scheme to discretize the first-order derivative (convection term). After the spatial discretization the resulting system of time dependent ODEs is stiff in nature [1].
@@ 4.3. Gray-Scott Model
$\nabla^{2}\phi_{i,j}\approx\frac{\phi_{i+1,j}+\phi_{i-1,j}+\phi_{i,j+1}+\phi_{i,j-1}-4\phi_{i,j}}{k^{2}}.$ (4.9)
@@
where $\Omega_{c}$ denotes a small square centered in $\Omega$. Although (4.8) is a well-known PDE for pattern formation, such patterns typically emerge only when the simulations are carried out over very long time intervals, e.g., $T=1000$ or $T=10000$. This is not feasible for the proposed schemes, since they are implicit in nature. Therefore, we refrain from demonstrating the pattern-formation phenomena.
@@
Figure 6. Convergence plot for Gray-Scott model
@@ 5. Results and Discussion
The Gray–Scott reaction–diffusion system provides a challenging test of robustness in PDEs with complex dynamics. Convergence and order are validated and are reported in Table 3 and Figure 6. The proposed integrators of order $p=1,3,$ produce less error than the $p=s$ class of [14], and the method of order $p=2,4,$ is competitive for the choice of parameters made while constructing the example method. A different choice of free parameter, along with a suitable minimization, would lead to a more efficient class with the same or better efficiency.
@@
  • Also, the order estimation given in the Figures 1, 3, 4 and 6 confirms that our proposed classes are almost free from any kind of order reduction. Moreover, our example methods produces small endpoint error norms than the method given in [14].
  • The numerical solution at $t=0.5$ for different diffusion coefficients, as illustrated in Figure 5 (5(a)-5(b)), preserves the parabolic character of the problem.
@@ 6. Conclusion
We have explained the construction of implicit GLMQS and presented four classes of $A$-stable and $L$-stable implicit numerical solvers for solving time-dependent stiff differential equations. The proposed classes of GLM are based on Nordsieck input vectors that satisfy the inherent quadratic stability criteria, along with the minimization of the error constant. Example methods of order $p=1,2,3,$ and $4$ are derived, and their governing coefficient matrices are reported in Section 3. We have also provided three numerical illustrations of time-dependent IVPs, both ODEs and PDEs, to test our proposed classes. The numerical results are reported for the van der Pol problem (ODE) and for both PDEs, the Burgers’ equation and the Gray-Scott model. These reported results show that our methods are convergent and achieve the expected order of accuracy. In future articles, we will be focusing on the construction of higher-order methods and the development of implicit-explicit numerical schemes using the present classes of implicit schemes for solving time-dependent partitioned differential equations.
Funding. The authors have received no funding.
Data availability. Not applicable.
Declarations. The authors report there are no competing interests to declare.
Acknowledgments. The authors are grateful to the editor and the anonymous reviewers for their valuable suggestions, which helped to improve the presentation of this paper.